Discounted Payback An investment project costs $13,500 and has annual cash flows of ars. What is the discounted payback period if the discount rate is 0 percent? What f the ount rate is 7 percent? If it is 21 percent? noturn Your firm is considering purchasing a machine with the following annual,

please show work
Answer #1

Answer::

given

Initial Investment P=$13500

Annual cash flow A=$3900

If discount rate r=0%

Discounted cashflow= A/(1+r)^n

where n= year in which cash flow occur

for year 1

Discounted cashflow D=3900/(1+0%)^1=$3900

Remaining value of  Investment=P-D=13500-3900=$9600

Since after 1 year we have to only recover $9600 so this become our investment for second year.

r 0%
Year Investment P Cashflow discounted cash flow D Remaining value of  Investment P-D
1 13500 3900 3900 9600
2 9600 3900 3900 5700
3 5700 3900 3900 1800
4 1800 3900 3900 -2100
5 -2100 3900 3900 -6000
6 -6000 3900 3900 -9900

Since payback period occur when Remaining value of  Investment =0

so from above table we found that payback period will occur between 3 and 4 years

So payback period =3+(1800/3900)=3.46 years

If discount rate r=7%

Discounted cashflow= A/(1+r)^n

where n= year in which cash flow occur

for year 1

Discounted cashflow D=3900/(1+7%)^1=$3644.86

Remaining value of  Investment=P-D=13500-3644.86=$9855.14

Since after 1 year we have to only recover $9844.15 so this become our investment for second year.

r 7%
Year Investment P Cashflow discounted cash flow D Remaining value of  Investment P-D
1 13500.00 3900.00 3644.86 9855.14
2 9855.14 3900.00 3406.41 6448.73
3 6448.73 3900.00 3183.56 3265.17
4 3265.17 3900.00 2975.29 289.88
5 289.88 3900.00 2780.65 -2490.77
6 -2490.77 3900.00 2598.73 -5089.50

Since payback period occur when Remaining value of  Investment =0

so from above table we found that payback period will occur between 4 and 5 years

So payback period =4+(289.88/2975.29)=4.097 years

If discount rate r=21%

Discounted cashflow= A/(1+r)^n

where n= year in which cash flow occur

for year 1

Discounted cashflow D=3900/(1+21%)^1=$3223.14

Remaining value of  Investment=P-D=13500-3644.86=$10276.86

Since after 1 year we have to only recover $10276.86 so this become our investment for second year.

r 21%
Year Investment P Cashflow discounted cash flow D Remaining value of  Investment P-D
1 13500.00 3900.00 3223.14 10276.86
2 10276.86 3900.00 2663.75 7613.11
3 7613.11 3900.00 2201.45 5411.66
4 5411.66 3900.00 1819.38 3592.28
5 3592.28 3900.00 1503.62 2088.66
6 2088.66 3900.00 1242.66 846.00

Since payback period occur when Remaining value of  Investment =0

So from above table we found that payback period will be more than 6 years